# Charming Darwin: the evolution of QCD parameters across different species

### v1

@@ 2.2 Systematic uncertainties
The size of systematic uncertainties in BESIII measurements of the differential spectra due to detector modelling are estimated by varying the assumed reconstruction resolution and the reconstruction and particle identification efficiencies, within their stated uncertainties and repeating the procedure described in Sec. 2.1. The calibration procedure described in Sec. 3 depends on the branching fractions that are assumed for the component exclusive branching fractions. Alternate spectra are produced based on the measured uncertainties of the exclusive branching fractions included in the fast simulation and checked to determine the magnitude of the dependence.
@@ 2.3 Background contributions
While we consider all foreseen detector effects on reconstructing and identifying signal events in the fast simulation, we do not consider possible contributions from backgrounds. Backgrounds will arise from the tag-side hadron being misidentified, the lepton candidate being a misidentified hadron, and contributions from $e^{+}e^{-}\gamma$ Dalitz decays of light mesons produced in the final state of charmed hadron decays. The first category can be handled in a similar fashion to the inclusive semileptonic beauty measurements at Belle and Belle II BelleQ2 ; BelleIIQ2 . The second category can be handled through a weighting procedure, where particle identification observables are used to weight the observed sample. The third and final category can be corrected for with a similar procedure to previous inclusive semileptonic charm measurements CLEO:2009uah ; BESIIIDsIncSL ; BESIII:2022cmg , where these backgrounds are corrected for with data control samples in which the identified lepton has opposite charge to the expected flavour of the parent hadron. The expected impact of background subtraction on the statistical precision are estimated from the purities observed in the previous inclusive semileptonic charm measurements CLEO:2009uah ; BESIIIDsIncSL ; BESIII:2022cmg , described in more detail in Sec. 5.
@@ 3 Calibration procedure and sensitivity of raw spectral moments
Figure 5: The linear calibration function determined for the first $q^{2}$ moment of the $D^{0}$ decay which is used to calibration the reconstructed moments to the corresponding generator moments.
@@
$C_{\mathrm{calib}}$ $=\langle q^{2n}_{\mathrm{gen,sel}}\rangle/\langle q^{2n}_{\mathrm{cal}}\rangle\qquad\text{or}\qquad\langle E^{n}_{\ell,\mathrm{gen,sel}}\rangle/\langle E^{n}_{\ell,\mathrm{cal}}\rangle\qquad,$ (2)
@@
$C_{\mathrm{gen}}$ $=\langle q^{2n}_{\mathrm{gen}}\rangle/\langle q^{2n}_{\mathrm{gen,sel}}\rangle\qquad\text{or}\qquad\langle E^{n}_{\ell,\mathrm{gen}}\rangle/\langle E^{2n}_{\ell,\mathrm{gen,sel}}\rangle\qquad$ (3)
@@
$\langle q^{2n}\rangle$ $=\sum_{i}q^{2n}_{i\,\text{cal}}\times C_{\mathrm{calib}}\times C_{\mathrm{gen}}\,,$ $\langle E_{\ell}^{n}\rangle$ $=\sum_{i}E^{n}_{\ell,i,\text{cal}}\times C_{\mathrm{calib}}\times C_{\mathrm{gen}}\,,$ (4)
@@ 4.1 Theoretical framework
The setup of the Operator Product Expansion (OPE) for inclusive $D\to X\ell\bar{\nu}_{\ell}$ decays differs from the standard, well-established setup for $B\to X_{c}\ell\bar{\nu}_{\ell}$ decays due to the hierarchy between the charm-quark and light-quark decay products. In the latter case, the ratio $m_{c}/m_{b}$ is fixed when setting up the OPE. However, for $D\to X_{s}$ and $D\to X_{d}$ decays, the $s$ and $d$ quarks need to be integrated out at their respective scales. This introduces four-quark operators (weak annihilation) (see e.g. Bigi:2009ym ; Breidenbach:2008ua for a detailed discussion). As such, the OPE used to describe charm decays resembles more that of the $B\to X_{u}\ell\bar{\nu}_{\ell}$ case. The setup of the HQE is such that the hadronic parameters depend on the mass of the initial state meson. Therefore, the hadronic matrix element differ between charm and beauty decays, but also between $D_{s}$ and $D$ decays.
@@
We note that our setup, which follows Fael:2019umf , differs from Gambino:2010jz ; Ligeti:2010vd ; Wu:2024jyf due to the three expansion parameters that we employ and through the definition of the four-quark operators given above. In the latter, the WA matrix elements are named $B_{\rm WA}$. Specifically, our expressions differ in the non-logarithmic terms from those of $b\to c$ due to a difference in the matching of the two-quark operators for the $c\to s$ transition. This subtle point was discussed in Fael:2019umf and overlooked in Gambino:2010jz ; Ligeti:2010vd ; Wu:2024jyf . In addition, converting $\tau_{0}$ to $B_{\rm WA}$ also requires setting the scale $\mu_{\rm WA}=0.8$ GeV.
The operators $T_{1}$ and $T_{2}$ discussed above are non-valence Weak Annihilation (WA) operators, which do not depend on the flavour of the spectator quark. For $D^{0}$ decays, only non-valence WA contributes, however for $D_{s}$ and $D^{+}$ decays the valence WA also enters weighted by the relevant CKM factors (see expressions in Sec. 6 of Fael:2019umf ). These operators are defined through (4.1) by replacing $s\to q$, where $q$ is the valence spectator quark. The non-valence WA contribution is similar for $D^{0}$ and $D^{+}$ but differs for $D_{s}$ decays due to $SU(3)$ breaking effects that also affect the other HQE elements. The valence and non-valence WA could be disentangled by measuring the WA parameters, $\tau_{0}$, from $D^{0},D^{+}$ and $D_{s}$ decays Bigi:2009ym . It is important to extract these WA from inclusive charm decays as they form a crucial input into the $D^{0},D^{+}$ and $D_{s}$ lifetime calculations.
@@
In the $B\to X_{c}\ell\bar{\nu}_{\ell}$ analyses, usually centralized moments are defined. For our exploratory study, we fit the raw moments as defined in (14), which can readily be converted to centralized moments. The expressions for the lepton-energy and $q^{2}$ moments at tree level are given in Fael:2019umf . For the $\Lambda_{c}$ decays, the expressions are similar to those for the $D$ decays, except that all $\sigma\cdot G$ terms cancel due to the spin structure and thus $\mu_{G}^{2}$ and $\rho_{LS}^{3}$ do not contribute [^4] Technically, we first have to split off the spatial and time derivatives in $\mu_{G}^{2}$ by converting from the RPI basis as $\mu_{G}^{2}\to(\mu_{G}^{2})^{\perp}-\rho_{D}^{3}/m_{c}-\rho_{LS}^{3}/m_{c}$. We can then set $(\mu_{G}^{2})^{\perp}$ and $\rho_{LS}^{3}$ to zero.. In the following, we consider moments up to $n=4$. We note that especially these higher order moments are very sensitive to the HQE parameters and would suffer from larger theoretical uncertainties due to neglected $1/m_{c}^{4}$ HQE parameters.
@@ 4.2 Theory challenges
For a precision determination of the HQE parameters in charm, the setup of the HQE for charm at leading order in $\alpha_{s}$ does not suffice due to the large value of the strong coupling constant at the relevant scale $m_{c}$. Practically this means higher-order $\alpha_{s}$ corrections should be included, while getting a reliable estimate for the missing higher-orders in $\alpha_{s}$ is challenging. Clearly, a simple scale variation (often done for $B$ decays) would render much larger uncertainties. For charm decays, $\alpha_{s}$ corrections for $q^{2}$ and lepton energy moments can be obtained from Fael:2022frj which uses an expansion in $\delta=1-\sqrt{\rho}$ with $\rho\equiv(m_{q}/m_{Q})^{2}$ for the $m_{Q}\to m_{q}$ transition. For massless quarks, the $\alpha_{s}$ corrections can be obtained from Chen:2022wit for $q^{2}$ moments. For the $\alpha_{s}^{2}$ BLM corrections are known for lepton energy moments Gambino:2010jz ; Aquila:2005hq (with lepton energy threshold cuts). For the total rate, $\alpha_{s}^{2}$ corrections are available Melnikov:2008qs . To our knowledge, for $q^{2}$ moments, no dedicated calculations for $c\to s$ transitions are available. They could be obtained by adapting existing calculations for inclusive $b\to c$ semileptonic transitions. For the latter, the full $\alpha_{s}^{2}$ corrections to the $q^{2}$ moments were recently obtained Fael:2024gyw  [^5] Slightly before that the BLM corrections for the $q^{2}$ moments were calculated Finauri:2023kte . For the lepton energy moments, non-BLM corrections are only known for $B$ decays for fixed $m_{c}/m_{b}$ masses and for several kinematical cuts Biswas:2009rb . In addition, for $b\to c$ decays even the $\alpha_{s}^{3}$ corrections to the total rate and kinematic moments without any kinematic cut are available Fael:2020tow ; Fael:2022frj .
In addition to the NNLO (or higher) corrections, a short-distance mass scheme for the charm quark is required to avoid the bad perturbative behavior of the pole mass. In $B\to X_{c}$ decays it is customary to use the kinetic mass Bigi:1994ga ; Bigi:1996si , which relies on a cut of $\Lambda_{\rm QCD}<\mu<m_{Q}$. For charm decays, this makes the allowed window for $\mu$ rather small. Refs. Gambino:2010jz ; Wu:2024jyf use the kinetic scheme for charm with $\mu=0.5$ GeV, however, as discussed in Fael:2018vsp (see also Ligeti:2010vd ), it remains to be seen if $\mu\simeq 0.5$ GeV is sufficiently away from $\Lambda_{\rm QCD}$. Alternatively, one could employ the $\overline{\rm MS}$ scheme at $\mu=m_{Q}$ Czarnecki:1994pu , although it has been argued that this scheme is less suitable for decays of the $b$ and $c$ quarks given that the typical energy release is much lower than $m_{Q}$ Melnikov:2000qh . In Ref. Ligeti:2010vd , the perturbative series of the $1S$ Hoang:1998ng ; Hoang:2005zw and PS scheme Beneke:1998rk , finding a reasonably well behaved series for the former.
@@
5 Towards an extraction of the Darwin operator and $|V_{cx}|$
To check the future potential of a inclusive charm program at BESIII, we perform an exploratory study and analyze the $q^{2}$ and $E_{\ell}$ moments. Given the theoretical challenges described above, we consider only the achievable experimental uncertainty on the extracted HQE parameters. Our sample data contains $\{D^{0},D^{+},D_{s},\Lambda_{c}\}\to(X_{s}+X_{d})\ell\bar{\nu}$ decays. As discussed, in the future these samples could be separated as is done for semileptonic $B\to X_{c}$ versus $B\to X_{u}$ decays. For this first study, we simply neglect the $c\to d$ contribution as it is suppressed by $|V_{cd}/V_{cs}|^{2}=0.04$. In addition, we correct for the lepton energy acceptance when applying the $\mathcal{C}_{\mathrm{gen}}$ factor, so that the measured moments have only the quoted $q^{2}_{th}$ threshold selections applied.
Specifically, we work up to order $1/m_{c}^{3}$ in the HQE and only include $\alpha_{s}$ corrections. It is customary to include an additional theoretical uncertainty due to not included higher orders in the HQE (and $\alpha_{s}$ expansion) (see e.g. the discussion in Finauri:2023kte ). We do not include any theory uncertainties in the fit and leave this for future studies. We also do not consider any external inputs for the $\mu_{G}^{2}$ parameter which can in principle be linked to the mass difference of the initial state and its first excited state.
For the strange quark mass, we use the $2+1+1$ Lattice QCD FLAG averages FlavourLatticeAveragingGroup:2019iem
@@
As discussed, the choice of the charm-quark mass definition is more challenging. Since our aim is to get an estimate for the attainable experimental precision only, we do not discuss in detail the different mass schemes for the quarks. For our study, we use the $\overline{\rm MS}$ definition, obtained from $2+1+1$ Lattice QCD FLAG average FlavourLatticeAveragingGroup:2019iem
@@
obtained from RunDec with $n_{f}=3$ active flavours Chetyrkin:2000yt . Recently, Wu:2024jyf used the inclusive charm branching ratios together with input for the HQE parameters from the $B\to X_{c}\ell\bar{\nu}_{\ell}$ decays to extract $\alpha_{s}$ employing the kinetic scheme for the charm mass. Our input value is in good agreement with the value extracted in that way. In a similar way as in Wu:2024jyf , a future moments analysis could also allow the determination of the value of the strong coupling constant directly from the data.
@@
depending only on the HQE parameters up to $\mathcal{O}(1/m_{b}^{3})$ and the weak annihilation coefficient $\tau_{0}$. The experimental data used in the fit are the fully inclusive ($q^{2}_{\rm th}=0\,\text{GeV}^{2}$) $\langle q^{2n}\rangle$ and $\langle E_{\ell}^{n}\rangle$ moments with $n=\{1,2,3,4\}$. We take into account the full statistical and systematic covariance between the two sets of spectral moments and the different orders. We perform the fit to the simulated data, construct an Asimov data set Cowan:2010js based on the fit result, and refit the Asimov data to determine the experimental precision. To take into account backgrounds from wrongly reconstructed tags, the statistical uncertainty is scaled by $1/\sqrt{\pi}$, where $\pi$ is the estimated purity based on the previous BESIII inclusive semileptonic analyses CLEO:2009uah ; BESIIIDsIncSL ; BESIII:2022cmg .
@@
The uncertainty budget of the $D^{0}$, $D^{+}$, $D_{s}$, and $\Lambda_{c}$ is given in Tables 4, 5, 6, and 7, respectively. In each decay mode, the statistical uncertainty can be improved with a larger BESIII data set, and the calibration uncertainty (labeled as MC stat.) can be improved with increases in the integrated luminosity of the simulated data set. The track reconstruction efficiency is also a leading uncertainty. We stress that the $D_{s}$ and $\Lambda_{c}$ modes specifically would benefit from more precise measurements of the exclusive branching ratios. We show the evolution of the precision estimates in Figure 9, where we assume that all uncertainties will scale with the square root of the integrated luminosities and that BESIII will keep 40 times the integrated luminosity for their simulation with respect to the recorded data. Increased samples of $D_{s}$ mesons and $\Lambda_{c}$ baryons at BESIII would allow for competitive precision on the HQE parameters for all four flavors of charmed hadrons.
@@
Full 21.35 29.13 1.39 37.47 66.12
Stat. 9.17 5.88 0.22 9.63 10.91
MC Stat. 14.69 16.37 0.30 22.64 27.08
$\epsilon_{\mathrm{track.}}$ 16.09 26.81 1.24 32.91 61.95
$\sigma_{\mathrm{track.}}$ 0.22 0.24 0.01 0.33 0.28
$\epsilon_{K_{S}^{0}}$ 1.79 0.77 0.01 1.60 0.94
$\sigma_{K_{S}^{0}}$ 0.04 0.03 0.00 0.03 0.05
$\epsilon_{\gamma}$ 0.72 0.81 0.06 1.13 1.69
$\sigma_{\gamma}$ 0.31 0.08 0.03 0.10 0.87
PID 0.03 0.13 0.00 0.07 0.14
$\mathcal{B}(D^{0}\to K_{1}^{-}\ell\nu_{\ell})$ 5.07 6.02 0.03 8.07 9.60
$\mathcal{B}(D^{0}\to K^{*-}\ell\nu_{\ell})$ 4.06 3.39 0.55 8.72 20.99
$\mathcal{B}(D^{0}\to K^{-}\ell\nu_{\ell})$ 3.64 4.50 0.21 5.38 4.95
$\mathcal{B}(D^{0}\to\rho^{-}\ell\nu_{\ell})$ 6.01 12.61 0.74 13.38 30.13
$\mathcal{B}(D^{0}\to(K\pi)_{S-\mathrm{wave}}\ell\nu_{\ell})$ 6.82 5.63 0.13 9.06 8.89

### v3

@@ 2.2 Systematic uncertainties
The size of systematic uncertainties in BESIII measurements of the differential spectra due to detector modeling are estimated by varying the assumed reconstruction resolution and the reconstruction and particle identification efficiencies within their stated uncertainties and repeating the procedure described in Sec. 2.1. The calibration procedure described in Sec. 3 depends on the branching fractions that are assumed for the component exclusive branching fractions. Alternate spectra are produced based on the measured uncertainties of the exclusive branching fractions included in the fast simulation and checked to determine the magnitude of the dependence.
The dominating experimental systematic uncertainties are associated with the measurements of the exclusive branching fractions, which are all currently limited by BESIII sample sizes, and corrections for imperfect simulation of detector response, which are all also dominated by the size of data control samples used to estimate the correction. As such, we assume that the experimental systematic uncertainties of each of the measurements will scale with the square root of the collected sample’s luminosity for data-taking scenarios possible in the near-future. We expect this assumption may not hold for samples larger than the currently collected $D^{0}$ and $D^{+}$ samples estimated in Table 1 or samples ten times larger than the listed $D^{+}_{s}$ and $\Lambda_{c}$ samples that could be collected at a proposed Super-Tau-Charm Factory STCF .
@@ 2.3 Background contributions
While we consider all foreseen detector effects on reconstructing and identifying signal events in the fast simulation, we do not consider possible contributions from backgrounds. Backgrounds will arise from the tag-side hadron being misidentified, the lepton candidate being a misidentified hadron, and contributions from $e^{+}e^{-}\gamma$ Dalitz decays of light mesons produced in the final state of charm hadron decays. The first category can be handled in a similar fashion to the inclusive semileptonic beauty measurements at Belle and Belle II BelleQ2 ; BelleIIQ2 . The second category can be handled through a weighting procedure, where particle identification observables are used to weight the observed sample. The third and final category can be corrected for with a similar procedure to previous inclusive semileptonic charm measurements CLEO:2009uah ; BESIIIDsIncSL ; BESIII:2022cmg , where these backgrounds are corrected for with data control samples in which the identified lepton has opposite charge to the expected flavour of the parent hadron. The expected impact of background subtraction on the statistical precision are estimated from the purities observed in the previous inclusive semileptonic charm measurements CLEO:2009uah ; BESIIIDsIncSL ; BESIII:2022cmg , described in more detail in Sec. 5.
@@ 3 Calibration procedure and sensitivity of raw spectral moments
Figure 5: The linear calibration function determined for the first $q^{2}$ moment of the $D^{0}$ decay which is used to calibrate the reconstructed moments to the corresponding generator moments.
@@
$\mathcal{C}_{\mathrm{calib}}$ $=\langle q^{2n}_{\mathrm{gen,sel}}\rangle/\langle q^{2n}_{\mathrm{cal}}\rangle\qquad\text{or}\qquad\langle E^{n}_{\ell,\mathrm{gen,sel}}\rangle/\langle E^{n}_{\ell,\mathrm{cal}}\rangle\qquad,$ (2)
@@
$\mathcal{C}_{\mathrm{gen}}$ $=\langle q^{2n}_{\mathrm{gen}}\rangle/\langle q^{2n}_{\mathrm{gen,sel}}\rangle\qquad\text{or}\qquad\langle E^{n}_{\ell,\mathrm{gen}}\rangle/\langle E^{2n}_{\ell,\mathrm{gen,sel}}\rangle\qquad$ (3)
@@
$\langle q^{2n}\rangle$ $=\sum_{i}q^{2n}_{i\,\text{cal}}\times\mathcal{C}_{\mathrm{calib}}\times\mathcal{C}_{\mathrm{gen}}\,,$ $\langle E_{\ell}^{n}\rangle$ $=\sum_{i}E^{n}_{\ell,i,\text{cal}}\times\mathcal{C}_{\mathrm{calib}}\times\mathcal{C}_{\mathrm{gen}}\,,$ (4)
@@ 4.1 Theoretical framework
The setup of the Operator Product Expansion (OPE) for inclusive $D\to X\ell\bar{\nu}_{\ell}$ decays differs from the standard, well-established setup for $B\to X_{c}\ell\bar{\nu}_{\ell}$ decays due to the hierarchy between the charm-quark and light-quark decay products. In the latter case, the ratio $m_{c}/m_{b}$ is fixed when setting up the OPE. However, for $D\to X_{s}$ and $D\to X_{d}$ decays, the $s$ and $d$ quarks need to be integrated out at their respective scales. This introduces four-quark operators (weak annihilation) (see e.g. Bigi:2009ym ; Breidenbach:2008ua for a detailed discussion). As such, the OPE used to describe charm decays resembles more that of the $B\to X_{u}\ell\bar{\nu}_{\ell}$ case. The setup of the HQE is such that the hadronic parameters depend on the mass of the initial state meson. Therefore, the hadronic matrix elements differ between charm and beauty decays, but also between $D_{s}$ and $D$ decays.
@@
We note that our setup, which follows Fael:2019umf , differs from Gambino:2010jz ; Ligeti:2010vd ; Wu:2024jyf due to the three expansion parameters that we employ and through the definition of the four-quark operators given above. In the latter, the WA matrix elements are named $B_{\rm WA}$. Specifically, our expressions differ in the non-logarithmic terms from those of $b\to c$ due to a difference in the matching of the two-quark operators for the $c\to s$ transition. This subtle point was discussed in Fael:2019umf and overlooked in Gambino:2010jz ; Ligeti:2010vd ; Wu:2024jyf . In addition, converting $\tau_{0}$ to $B_{\rm WA}$ also requires setting the scale $\mu_{\rm WA}$, which in Gambino:2010jz was set at $\mu_{\rm WA}=0.8$ GeV.
The operators $T_{1}$ and $T_{2}$ discussed above are non-valence Weak Annihilation (WA) operators, which do not depend on the flavour of the spectator quark. For $D^{0}$ decays, only non-valence WA contributes, however for $D_{s}$ and $D^{+}$ decays the valence WA also enters weighted by the relevant CKM factors (see expressions in Sec. 6 of Fael:2019umf ). These operators are defined through (4.1) by replacing $s\to q$, where $q$ is the valence spectator quark. The non-valence WA contribution is similar for $D^{0}$ and $D^{+}$ but differs for $D_{s}$ decays due to $SU(3)$ breaking effects that also affect the other HQE elements. The valence and non-valence WA could be disentangled by measuring the WA parameters, $\tau_{0}$, from $D^{0},D^{+}$ and $D_{s}$ decays Bigi:2009ym . It is important to extract these WA parameters from inclusive charm decays as they form a crucial input into the $D^{0},D^{+}$ and $D_{s}$ lifetime calculations.
@@
These moments can be converted to centralized moments, which are customary to use $B\to X_{c}\ell\bar{\nu}_{\ell}$ analyses (see e.g. Finauri:2023kte ; Bernlochner:2022ucr ). For our exploratory study, we however fit the raw moments as defined in (14). The expressions for the lepton-energy and $q^{2}$ moments at tree level are given in Fael:2019umf . For the $\Lambda_{c}$ decays, the expressions are similar to those for the $D$ decays, except that all $\sigma\cdot G$ terms cancel due to the spin structure and thus $\mu_{G}^{2}$ and $\rho_{LS}^{3}$ do not contribute [^4] Technically, we first have to split off the spatial and time derivatives in $\mu_{G}^{2}$ by converting from the RPI basis as $\mu_{G}^{2}\to(\mu_{G}^{2})^{\perp}-\rho_{D}^{3}/m_{c}-\rho_{LS}^{3}/m_{c}$. We can then set $(\mu_{G}^{2})^{\perp}$ and $\rho_{LS}^{3}$ to zero.. In the following, we consider moments up to $n=4$. We note that especially these higher order moments are very sensitive to the HQE parameters and would suffer from larger theoretical uncertainties due to neglected $1/m_{c}^{4}$ HQE parameters.
@@ 4.2 Theory challenges
For a precision determination of the HQE parameters in charm, the setup of the HQE for charm at leading order in $\alpha_{s}$ does not suffice due to the large value of the strong coupling constant at the relevant scale $m_{c}$. Practically this means higher-order $\alpha_{s}$ corrections should be included, while getting a reliable estimate for the missing higher-orders in $\alpha_{s}$ is challenging. Clearly, a simple scale variation (often done for $B$ decays) would render much larger uncertainties. For charm decays, $\alpha_{s}$ corrections for $q^{2}$ and lepton energy moments can be obtained from Fael:2022frj which uses an expansion in $\delta=1-\sqrt{\rho}$ with $\rho\equiv(m_{q}/m_{Q})^{2}$ for the $m_{Q}\to m_{q}$ transition. For massless quarks, the $\alpha_{s}$ corrections can be obtained from Chen:2022wit for $q^{2}$ moments. For the lepton energy moments in charm decays, Gambino:2010jz implemented the $\alpha_{s}^{2}$ BLM corrections obtained from Aquila:2005hq for different lepton energy threshold cuts. For the total rate, $\alpha_{s}^{2}$ corrections are available Melnikov:2008qs . To our knowledge, for $q^{2}$ moments, no dedicated calculations for $c\to s$ transitions are available. They could be obtained by adapting existing calculations for inclusive $b\to c$ semileptonic transitions. For the latter, the full $\alpha_{s}^{2}$ corrections to the $q^{2}$ moments were recently obtained Fael:2024gyw  [^5] Slightly before that the BLM corrections for the $q^{2}$ moments were calculated Finauri:2023kte . For the lepton energy moments, non-BLM corrections are only known for $B$ decays for fixed $m_{c}/m_{b}$ masses and for several kinematical cuts Biswas:2009rb . In addition, for $b\to c$ decays even the $\alpha_{s}^{3}$ corrections to the total rate and kinematic moments without any kinematic cut are available Fael:2020tow ; Fael:2022frj .
In addition to the NNLO (or higher) corrections, a short-distance mass scheme for the charm quark is required to avoid the bad perturbative behavior of the pole mass. In $B\to X_{c}$ decays it is customary to use the kinetic mass Bigi:1994ga ; Bigi:1996si , which relies on a cut of $\Lambda_{\rm QCD}<\mu<m_{Q}$. See also Refs. Bauer:2004ve ; Bauer:2002sh for a discussion on different mass schemes in $B$ decays. For charm decays, this makes the allowed window for $\mu$ rather small. Refs. Gambino:2010jz ; Wu:2024jyf use the kinetic scheme for charm with $\mu=0.5$ GeV, however, as discussed in Fael:2018vsp (see also Ligeti:2010vd ), it remains to be seen if $\mu\simeq 0.5$ GeV is sufficiently away from $\Lambda_{\rm QCD}$. Alternatively, one could employ the $\overline{\rm MS}$ scheme at $\mu=m_{Q}$ Czarnecki:1994pu , although it has been argued that this scheme is less suitable for decays of the $b$ and $c$ quarks given that the typical energy release is much lower than $m_{Q}$ Melnikov:2000qh . In Ref. Ligeti:2010vd , the perturbative series of the $1S$ Hoang:1998ng ; Hoang:2005zw and PS scheme Beneke:1998rk , finding a reasonably well behaved series for the former.
@@
5 Towards an extraction of the Darwin and other HQE operators
To check the future potential of an inclusive charm program at BESIII, we perform an exploratory study and analyze the $q^{2}$ and $E_{\ell}$ moments. Given the theoretical challenges described above, we consider only the achievable experimental uncertainty on the extracted HQE parameters. Our sample data contains $\{D^{0},D^{+},D_{s},\Lambda_{c}\}\to(X_{s}+X_{d})\ell\bar{\nu}$ decays. As discussed, in the future these samples could be separated as is done for semileptonic $B\to X_{c}$ versus $B\to X_{u}$ decays. For this first study, we neglect the $c\to d$ contribution in the theoretical expressions for simplicity. In addition, we correct for the lepton energy acceptance when applying the $\mathcal{C}_{\mathrm{gen}}$ factor, so that the measured moments have only the quoted $q^{2}_{th}$ threshold selections applied.
Specifically, we work up to order $1/m_{c}^{3}$ in the HQE and only include $\alpha_{s}$ corrections. It is customary to include an additional theoretical uncertainty due to excluded higher orders in the HQE (and $\alpha_{s}$ expansion) (see e.g. the discussion in Finauri:2023kte ). We do not include any theory uncertainties in the fit, as the aim of our work is to determine the experimental feasibility of such measurements. Estimates of the theory uncertainties requires a dedicated program addressing the challenges detailed above We also do not consider any external inputs for the $\mu_{G}^{2}$ parameter which can in principle be linked to the mass difference of the initial state and its first excited state.
For the strange quark mass, we use the $2+1+1$ Lattice QCD FLAG averages FlavourLatticeAveragingGroup:2019iem ; EuropeanTwistedMass:2014osg ; FermilabLattice:2018est ; Chakraborty:2014aca ; Lytle:2018evc
@@ 5 Towards an extraction of the Darwin and other HQE operators
As discussed, the choice of the charm-quark mass definition is more challenging. Since our aim is to get an estimate for the attainable experimental precision only, we do not discuss in detail the different mass schemes for the quarks. For our study, we use the $\overline{\rm MS}$ definition, obtained from $2+1+1$ Lattice QCD FLAG average FlavourLatticeAveragingGroup:2019iem ; EuropeanTwistedMass:2014osg ; FermilabLattice:2018est ; Chakraborty:2014aca ; Lytle:2018evc ; Alexandrou:2014sha
@@
obtained from RunDec with $n_{f}=3$ active flavours Chetyrkin:2000yt ; Schmidt:2012az ; Herren:2017osy . Recently, Wu:2024jyf used the inclusive charm branching ratios together with input for the HQE parameters from the $B\to X_{c}\ell\bar{\nu}_{\ell}$ decays to extract $\alpha_{s}$ employing the kinetic scheme for the charm mass. Our input value is in good agreement with the value extracted in that way. In a similar way as in Wu:2024jyf , a future moments analysis could also allow the determination of the value of the strong coupling constant directly from the data.
@@
depending only on the HQE parameters up to $\mathcal{O}(1/m_{b}^{3})$ and the weak annihilation coefficient $\tau_{0}$. The experimental data used in the fit are the fully inclusive ($q^{2}_{\rm th}=0\,\text{GeV}^{2}$) $\langle q^{2n}\rangle$ and $\langle E_{\ell}^{n}\rangle$ moments with $n=\{1,2,3,4\}$. We take into account the full statistical and systematic covariance between the two sets of spectral moments and the different orders. We perform the fit to the simulated data, construct an Asimov data set Cowan:2010js based on the fit result, and refit the Asimov data to determine the experimental precision. To take into account backgrounds from wrongly reconstructed tags, the statistical uncertainty is scaled by $1/\sqrt{P}$, where $P$ is the estimated purity based on the previous BESIII inclusive semileptonic analyses CLEO:2009uah ; BESIIIDsIncSL ; BESIII:2022cmg .
@@
The uncertainty budget of the $D^{0}$, $D^{+}$, $D_{s}$, and $\Lambda_{c}$ is given in Tables 4, 5, 6, and 7, respectively. In each decay mode, the statistical uncertainty can be improved with a larger BESIII data set, and the calibration uncertainty (labeled as ‘MC stat.’) can be improved with increases in the integrated luminosity of the simulated data set. The track reconstruction efficiency is also a leading uncertainty. We stress that the $D_{s}$ and $\Lambda_{c}$ modes specifically would benefit from more precise measurements of the exclusive branching ratios. We show the evolution of the precision estimates in Figure 9, where we assume that all uncertainties will scale with the square root of the integrated luminosities and that BESIII will keep 40 times the integrated luminosity for their simulation with respect to the recorded data. Increased samples of $D_{s}$ mesons and $\Lambda_{c}$ baryons at BESIII would allow for competitive precision on the HQE parameters for all four flavours of charm hadrons.
@@
Full 21.4 29.1 1.4 37.5 66.1
Stat. 9.2 5.9 0.2 9.6 10.9
MC Stat. 14.7 16.4 0.3 22.6 27.1
$\epsilon_{\mathrm{track.}}$ 16.1 26.8 1.2 32.9 62.0
$\sigma_{\mathrm{track.}}$ 0.2 0.2 0.0 0.3 0.3
$\epsilon_{K_{S}^{0}}$ 1.8 0.8 0.0 1.6 0.9
$\sigma_{K_{S}^{0}}$ 0.0 0.0 0.0 0.0 0.1
$\epsilon_{\gamma}$ 0.7 0.8 0.1 1.1 1.7
$\sigma_{\gamma}$ 0.3 0.1 0.0 0.1 0.9
PID 0.0 0.1 0.0 0.1 0.1
$\mathcal{B}(D^{0}\to K_{1}^{-}\ell\nu_{\ell})$ 5.1 6.0 0.0 8.1 9.6
$\mathcal{B}(D^{0}\to K^{*-}\ell\nu_{\ell})$ 4.1 3.4 0.6 8.7 21.0
$\mathcal{B}(D^{0}\to K^{-}\ell\nu_{\ell})$ 3.6 4.5 0.2 5.4 5.0
$\mathcal{B}(D^{0}\to\rho^{-}\ell\nu_{\ell})$ 6.0 12.6 0.7 13.4 30.1
$\mathcal{B}(D^{0}\to(K\pi)_{S-\mathrm{wave}}\ell\nu_{\ell})$ 6.8 5.6 0.1 9.1 8.9